Eva is 10 years older than Feni. In 8 years, twice Eva's age will equal to three times Feni's age. How old is each at present.
Answer by Maths68(1474)

Eva is 10 years older than Feni. In 8 years, twice Eva's age will equal to three times Feni's age. How old is each at present
Let
Present age of Eva = e years old
Present age of Feni = f years old


Eva is 10 years older than Feni.
e=f+10..............(1)
In 8 years,
Eva will be = e+8
Feni will be = f+8
Then
Twice Eva's age will equal to three times Feni's age
2(e+8)=3(f+8)
2e+16=3f+24
2e=3f+24-16
2e=3f+8..............(2)
Substitute the value of e from (1) to (2)
2e=3f+8
2(f+10)=3f+8
2f+20=3f+8
2f-3f=8-20
-f=-12
Multiply by -1 both sides
f=12
Put the value of f in (1)
e=f+10..............(1)
e=12+10
e=22


Present age of Eva = e = 22 years old
Present age of Feni = f = 12 years old